# A Lesson in Splats: Teacher-Guided Diffusion for 3D Gaussian Splats Generation with 2D Supervision

### v2

@@ 4.1 Experimental Setups
(a.1) Splatter Image 24.1992 0.9213 0.0843 31.1158 0.9808 0.0269
(a.2) feedforwad (our architecture) 19.9947 0.8613 0.1588 23.2363 0.9165 0.0955
@@
(c.1) stage II w timestep from $[0,T]$ 24.6998 0.9233 0.0907 33.1196 0.9820 0.0205
(c.2) stage II w timestep from $[t^{*},T]$ 24.8324 0.9300 0.0871 33.4658 0.9883 0.0176
(d.1) stage I w/o consistency 22.6078 0.9046 0.1083 28.2025 0.9690 0.0411
(d.2) stage I w consistency 23.7293 0.9181 0.0979 29.9227 0.9774 0.0254
(d.3) stage I, stage II w consistency 24.9137 0.9332 0.0847 33.7061 0.9886 0.0153
Table 4: Ablations Studies on Single view Reconstruction, evaluated on the validation set of ShapeNet-SRN Cars
@@ 4.3 Image conditioned reconstruction
ShapeNet-SRN. We benchmark our diffusion model on the ShapeNet-SRN dataset, as presented in Tab. 1. Using only a single input view, our model achieves PSNR improvements of $0.84$ and $0.78$ on the cars and chairs splits, respectively, compared to the Splatter Image baseline.
For qualitative evaluation, we compare our method with Splatter Image, which serves as our teacher model in Fig. 1 and in Fig. 3. As seen in the first row (Fig. 3 (a)), images generated by Splatter Image occasionally exhibit artifacts and distortions. In contrast, our model generally produces more fine-grained structures and higher-quality details. Furthermore, as shown in Fig.  reffig:qual (b), the Gaussians generated by our model are denser and exhibit regular shapes, whereas those produced by Splatter Image tend to be oversized and less uniform.
RealEstate10K. We evaluate our method against recent state-of-the-art approaches on the real-world RealEstate10K dataset. As shown in Tab. 3, our model outperforms the teacher network, Flash3D, across three different evaluation settings, achieving an average PSNR improvement of $0.5$. The visual comparisons in Fig. 1 and Fig. 3 further demonstrates the superiority of our method, consistently producing cleaner images while Flash3D struggles in unseen regions, resulting in blurry artifacts.
@@ 4.4 Additional View Guidance
Table 5 compares our view guidance method to a 2-view 3DGS optimization procedure, as outlined by [21], which is initialized using the base model. Our diffusion model demonstrates a $0.2$ PSNR improvement when incorporating image guidance, with an additional $0.2$ PSNR gain achieved through Gaussian Splats optimization, consistently outperforming the Splatter Image baseline. While here we demonstrate guidance in a two-view settings, the guidance mechanism can naturally be extended to multiview scenarios, which we leave for future exploration.
@@
Figure 3: Qualitative results. (a) Qualitative comparison on the ShapeNet-SRN dataset for the Car and Chair categories. Our model produces views that are more faithful to the source image and better maintain plausibility. (b) Comparison of Gaussian Splat outputs between Splatter Image and our diffusion model shows that our model generates more regular patterns that closely follow the object surface. (c) Scene-level qualitative comparison on the RealEstate10K dataset demonstrates that our method produces more realistic results, particularly in ambiguous areas, such as the 2D edge separating the bed and the floor.
@@ 4.5 Ablation
Architectural Comparison. To assess whether our improvements stem from the diffusion framework or architectural changes, we trained a feedforward model using the same U-Net architecture as our diffusion model, with the timestep-related layers removed to create a deterministic version. This feedforward network directly predict gaussian parameter from the input image as Splatter Image. Due to a much smaller model size, it performs significantly worse than Splatter Image (Tab. 4.a), underscoring that the diffusion approach is essential to the enhanced results we observe.
@@
Cycle consistency. By introducing a feedback loop in which the predicted target view images are rendered back to the source view and supervised with the ground truth input image, we achieve performance improvements in both stages, as demonstrated in Tab. 4(d).
@@ Acknowledgment
Or Litany is a Taub fellow and is supported by the Azrieli Foundation Early Career Faculty Fellowship. This research was supported in part by an academic gift from Meta.
@@ 6 Additional results
6.1 Additional qualitative results
We present visual comparisons of our method to PixelNeRF [61] and VisionNeRF [27] on ShapeNet-SRN Cars and Chairs in Fig. 5. More qualitative results from RealEstate10K dataset is in Fig. 6.
Figure 5: Additional qualitative results. Qualitative comparisons on the ShapeNet-SRN dataset for additional viewpoints and objects from the Car and Chair categories. Our model produces views that are more faithful to the source image and better maintain plausibility, while maintaining the fast rendering of Splatter Image.
6.2 Additional ablations
@@ 7.1 ShapeNet-SRN Cars and Chairs
We adhere to the standard protocol for the ShapeNet-SRN dataset. Specifically, we use the provided images, camera intrinsics, camera poses, and data splits provided by [44] with a resolution of $128\times 128$. Our method is trained using relative camera poses. For single-view reconstruction, view 64 serves as the conditioning view, while for additional-view guidance, views 88 is used as guidance view. All remaining available views are treated as target views, where we compute novel view synthesis metrics.
@@ 7.2 RealEstate10K
We obtain 65,384 videos and their corresponding camera pose trajectories from the provided youtube links. Using these camera poses, we perform sparse point cloud reconstruction with COLMAP [41]. For evaluation, we adopt the test split provided by MINE [26] and follow prior work by assessing PSNR on novel frames that are 5 and 10 frames ahead of the source frame. Additionally, we evaluate on a randomly sampled frame within an interval of $\pm$30 frames, using the same frames employed in MINE’s evaluation. For evaluation, we use a total of 3,205 frames. The results presented in Tab. 3 are sourced from Flash3D [47]. Our model is trained and tested at a resolution of $256\times 384$.
@@ 8 Implementation details
Multi-step Denoising. We train the model on 4 NVIDIA A6000 GPUs. Our diffusion model is quite efficient . For bootstrapping at stage 1, we use a batch size of $100$ on each GPU. After obtaining the diffusion model from the teacher model, we perform multi-step denoising with a DDIM sampler of 10 inference steps. The batch size for stage 2 reduces to 10. We assign different weights to the rendering loss obtained at different timesteps and accumulate them for back-propagation throughout the denoising steps.
@@
Method GS optim Guidance PSNR SSIM LPIPS
Splatter Image 24.75 0.93 0.06
25.24 0.94 0.06
Ours 25.18 0.93 0.06
25.26 0.94 0.06
25.36 0.94 0.06
25.55 0.95 0.05

### v3

@@ 4.2 Implementation Details
(a.1) Feedforward, Splatter Image (Large) 24.1992 0.9213 0.0843 31.1158 0.9808 0.0269
(a.2) Feedforward, Splatter Image (Medium) 19.9947 0.8613 0.1588 23.2363 0.9165 0.0955
(a.3) Our diffusion (Medium), Medium teacher model 21.7506 0.8910 0.1093 28.0276 0.9621 0.0452
@@
(c.1) stage I w/o consistency 22.6078 0.9046 0.1083 28.2025 0.9690 0.0411
(c.1) stage I w consistency 23.7293 0.9181 0.0979 29.9227 0.9774 0.0254
(c.3) stage I w, stage II w/o consistency 24.6897 0.9229 0.0912 33.0582 0.9805 0.0211
(c.4) stage I, stage II w consistency (full model) 24.9137 0.9332 0.0847 33.7061 0.9886 0.0153
Table 4: Ablations Studies on Single view Reconstruction, evaluated on the validation set of ShapeNet-SRN Cars. In (b) and (c) rows, we use Splatter Image (Large) as a teacher to train our diffusion model (Medium).
@@ 4.3 Image Conditioned Reconstruction
ShapeNet-SRN. We benchmark our diffusion model on the ShapeNet-SRN dataset, as presented in Tab. 1. Using only a single input view, our model achieves PSNR improvements of 0.840.840.84 and 0.780.780.78 on the cars and chairs splits, respectively, compared to the Splatter Image baseline.
For qualitative evaluation, we compare our method with Splatter Image, which serves as our teacher model in Fig. 1 and in Fig. 3. As seen in the first row (Fig. 3 (a)), images generated by Splatter Image occasionally exhibit artifacts and distortions. In contrast, our model generally produces more fine-grained geometric structures and higher-quality details. Furthermore, as shown in Fig.  3 (b), the Gaussians generated by our model are denser and exhibit regular shapes, whereas those produced by Splatter Image tend to be oversized and less uniform.
RealEstate10K. We evaluate our method against recent state-of-the-art approaches on the real-world RealEstate10K dataset. As shown in Tab. 2, our model outperforms the teacher network, Flash3D, across three different evaluation settings, achieving an average PSNR improvement of 0.50.50.5. The visual comparisons in Fig. 1 and Fig. 3 further demonstrates the superiority of our method, consistently producing cleaner images while Flash3D struggles in unseen regions, resulting in blurry artifacts.
@@ 4.4 Additional View Guidance
Table 5 compares our view guidance method to a 2-view 3DGS optimization procedure, as outlined by [23], which is initialized using the base model. Our diffusion model demonstrates a 0.10.10.1 PSNR improvement for using 3D GS optimization and a 0.20.20.2 PSNR improvement when incorporating image guidance, with an additional 0.20.20.2 PSNR gain achieved through Gaussian Splats optimization, consistently outperforming the Splatter Image baseline, where guidance is not feasible. While here we demonstrate guidance in a two-view settings, the guidance mechanism can naturally be extended to multiview scenarios.
@@
Figure 3: Qualitative results. (a) Qualitative comparison on the ShapeNet-SRN dataset. Our model produces views that are more faithful to the source image and better maintain plausibility. (b) Comparison of Gaussian Splat outputs between Splatter Image and our diffusion model shows that our model generates more regular patterns that closely follow the object surface. (c) Scene-level qualitative comparison on the RealEstate10K dataset demonstrates that our method produces more realistic results, particularly in ambiguous areas, such as the 2D edge separating the bed and the floor. "M" and "L" denote "Medium" and "Large".
@@ 4.5 Ablation
Architectural Comparison. Our diffusion model uses a smaller U-Net architecture (Medium) than the original Splatter Image (Large) (Tab.4 (a.1)). To assess whether our improvements stem from the diffusion framework or architectural changes, we trained a feedforward model the same size as our diffusion model, which we refer to as Splatter Image (Medium) (Tab.4 (a.2)). Due to a smaller model size, it performs significantly worse than Splatter Image (Large) as seen in Fig. 3. With both medium and large teacher models, our diffusion model can significantly enhance the results of the base teacher.
@@
Cycle consistency. By introducing a feedback loop in which the predicted target view images are rendered back to the source view and supervised with the ground truth input image, we achieve performance improvements in both stages, as demonstrated in Tab. 4(c).
@@ Acknowledgment
Or Litany is a Taub fellow and is supported by the Azrieli Foundation Early Career Faculty Fellowship. He is also supported by the Israel Science Foundation through a personal grant (ISF 624/25) and an equipment grant (ISF 2903/25). This research was supported in part by an academic gift from Meta.
@@ 6 Additional results
6.1 Quantitative results
Co3D is an object-level dataset captured in the real world. We train our model on the Co3D hydrant class (with Splatter Image [55] as the teacher model) and compared it against ViewSet Diffusion [53] and Splatter Image [55] in Table 6. For both ViewSet and our model, we report average scores over seeds and across the test data.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
ViewSet Diffusion [53] 21.24 0.79 0.201
Splatter Image [55] 21.77 0.78 0.154
Ours 22.34 0.82 0.149
Table 6: Comparison on Co3D hydrant dataset.
6.2 Qualitative results
We present visual comparisons of our method to PixelNeRF [68] and VisionNeRF [28] on ShapeNet-SRN Cars and Chairs in Fig. 5. Although our diffusion model is of smaller size (Medium) than the original Splatter Image (Large), we are still able to outperform it. More qualitative results from RealEstate10K dataset are in Fig. 6.
Figure 5: Additional qualitative results. Qualitative comparisons on the ShapeNet-SRN dataset for additional viewpoints and objects from the Car and Chair categories. Our model produces views that are more faithful to the source image and better maintain plausibility, while maintaining the fast rendering of Splatter Image. Note that while our diffusion model is of smaller size (Medium) than the original Splatter Image (Large), we are still able to outperform it.
6.3 Ablations
@@ 7.1 ShapeNet-SRN Cars and Chairs
We adhere to the standard protocol for the ShapeNet-SRN dataset. Specifically, we use the provided images, camera intrinsics, camera poses, and data splits provided by [51] with a resolution of $\times 128$ 128. Our method is trained using relative camera poses. For single-view reconstruction, view 64 serves as the conditioning view, while for additional-view guidance, views 88 is used as guidance view. All remaining available views are treated as target views, where we compute novel view synthesis metrics.
@@ 7.2 RealEstate10K
We obtain 65,384 videos and their corresponding camera pose trajectories from the provided youtube links. Using these camera poses, we perform sparse point cloud reconstruction with COLMAP [48]. For evaluation, we adopt the test split provided by MINE [27] and follow prior work by assessing PSNR on novel frames that are 5 and 10 frames ahead of the source frame. Additionally, we evaluate on a randomly sampled frame within an interval of $\pm$30 frames, using the same frames employed in MINE’s evaluation. For evaluation, we use a total of 3,205 frames. The results presented in Tab. 2 are sourced from Flash3D [54]. Our model is trained and tested at a resolution of 256×384256$\times$ 384256 × 384.
@@ 8 Implementation details
Multi-step Denoising. We train the model on 4 NVIDIA A6000 GPUs. Our diffusion model is quite efficient . For bootstrapping at stage 1, we use a batch size of 100100100 on each GPU. After obtaining the diffusion model from the teacher model, we perform multi-step denoising with a DDIM sampler of 10 inference steps. The batch size for stage 2 reduces to 10. We assign different weights to the rendering loss obtained at different timesteps and accumulate them for back-propagation throughout the denoising steps.
Misc. We use Adam [24] as our optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$. We use a total noising steps of 100100100, with a linear scheduling, starting from 0.00010.00010.0001 to 0.20.20.2. For the bootstrapping stage, we use the teacher model to provide both supervision and noised samples. Instead of predicting the noises added to the splatters, our diffusion model denoises the noised inputs to clean samples directly. We trained for 5000 epochs for the bootstrapping stage and then finetune for 1000 epochs. For a fair comparison, we further finetuned the Splatter Image model for 1000 epochs and found negligible improvement because the model has converged. We set $t^{*}$ to be 20. For the architecture of diffusion model, we use the U-Net implementation from diffusers  [^1] https://huggingface.co/docs/diffusers. For the consistency branch, we use a denoising step of 10.
Method GS optim Guidance PSNR SSIM LPIPS
Splatter Image 24.75 0.93 0.06
25.24 0.94 0.06
Ours 25.18 0.93 0.06
25.26 0.94 0.06
25.36 0.94 0.06
25.55 0.95 0.05